OpenStudy (anna_13):

Max observes the zoo and the library from a helicopter flying at a height of 200 times square root of 3 feet above the ground, as shown below: A helicopter is flying at a height of 200 multiplied by square root of 3 feet above the ground. A zoo and a library are on the ground on the same side of the helicopter. The angle made by the line joining the helicopter and the zoo with the ground is 60 degrees. The angle made by the line joining the helicopter and the library with the ground is 30 degrees. What is the distance between the zoo and the library?

OpenStudy (anna_13):

@freckles

OpenStudy (anonymous):

\[200 \sqrt{2} \{\cot (30 {}^{\circ})-\cot (60 {}^{\circ})\}=400 \sqrt{\frac{2}{3}}=326.599 \]